UT Nov 05 2022 - Split LBTI Science (1st half) / DEC (2nd half)




Night Info

  • Sunset 17:26
  • Nautical Twilight Ends 18:21
  • Astro Twilight Ends 18:50
  • Night Center 00:03
  • Switch from LBTI to bino-LUCI ~7min - Handover 00:00
  • Astro Twilight Begins 05:17
  • Nautical Twilight Begins 05:46
  • Sunrise 06:41
  • Moonset 03:54
  • Moon illumination ~86%

Staff

  • OSA: D.G.Huerta
  • ISp: Mtn
  • Mtn Manager: B.Bingham
  • ISA/SO: B.Rothberg, R.T. Gatto (DEC time)
  • AO: D. Miller

Night Plan

ALTA prediction for Tonight

20221104_see_evol_time_Height_20_20000.png

Starting Config LBTI Science

LUCI AO/ESM

  • Files:
    • Service_Observing/2022B_D_E/LUCI/AO/AO_Testing_V105.xml
      • REQUIRES NEW VERSION OF OT if any sub-frame scripts are generated
    • Service_Observing/2022B_D_E/LUCI/ESM/ESM_V46.xml
      • for mixed mode ESM-SL scripts have already been generated and edited
  • Thin clouds, seeing 1.2" or better for AO - more flexible for ESM
DONE OSCO
ESM On-Sky Checks
  • ESM Gain Tests
    • Check each Bin - adjust gains - measure IQ in field
    • /home/lbto/Service_Observing/2022B_D_E/LUCI/ESM/
      • DONE /M52 Bin 1 (R=10.6) - twilight-02:30hr
        • use narrow band filters to measure PSF of AO Ref star
      • DONE /M52 Bin 4 (R=15.6) - twilight-02:30hr
      • /M34 Bin 2 (R=12.4) - all night
        • use narrow band filters to measure PSF of AO Ref star
  • DONE NS Target Didymos - would be done before the AO version - 0.5hr
    • Push limits of faint Bin 4 - use clear OCAM filter
    • Visible 01:30-twilight
D-Time AO Observations
  • AO/ESM Observations of JWST PSF stars
      • OT: Service_Observing/AZ_queue/AZMini-Q2022B/Poor Weather Programs/LD-2022B-001_Yang
      • Goal is to constrain whether stars are binary or single down to the limits of AO/ESM - can also be used as a check of the system - about 30min w/overheads per target
    • DONE Target 4 RA 10hr Dec +21 -visible 03:15-twilight - Full AO - Bin 2 - Highest Priority
    • DONE Target 1 RA 03hr Dec -18 - visible 22:30-03hr - optimal 00:45 Full AO - Bin 3 or ESM Bin 4 - if conditions are very good, use Full AO, conditions less good use ESM - 2nd priority
    • DONE Target 2 RA 03hr Dec -18 - visible 22:30-03hr - optimal 00:45 Full AO - Bin 2 - 2nd priority
    • Target 3 RA 10hr Dec +21 -visible 03:15-twilight - only execute in the best conditions - lowest priority
ESM On-Sky Checks
  • DONE Arp 187 - 0.75 hr
    • Full AO LUCI-1, SL LUCI-2
    • Bin 2/3 Test - IQ w/Full AO+N375 camera
    • also tests bayside stages
      • Service_Observing/2022B_D_E/LUCI/AO/AO_Testing/Arp187_AO_N375_Bino.xml
      • Service_Observing/2022B_D_E/LUCI/AO/AO_Testing/Arp187_AO_N375_L1.xml

Poor Weather Backup

  • Photometric Standards (LUCI) - only if clear - any seeing
    • /home/lbto/Service_Observing/2022B_D_E/PoorWeather/LUCI_Phot_ZP_OT_V6.xml
  • Local Mergers - bino LUCI
    • /home/lbto/Service_Observing/2022B_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V54.xml
      • Finder Charts: /home/lbto/Service_Observing/2022B_D_E/PoorWeather/Rothberg/LocalMergers/finder_charts/
        • jpgs of FOV matched by name
      • NGC 1210 - RA 03hr Dec -25 - visible 23:30-01:30 - elevation never gets above 33
      • NGC 2744 - RA 09h4 Dec +18 - visible 02:15-twilight
      • Arp 230 - RA 00 Dec -13 - visible 19:30-01hr
        • NOTE: there are two tellurics - use late telluric for 2nd half of night
        • Also a MODS script to run if nothing else to do


Night Log

Closed Dome

  • UT 00:17 - logged into OBS2 - too many sessions on OBS1 - OBS2 had a leftover session with ARGOS Swingarm GUIs and an old VNC session. Killed those and faster response
  • SX BCU2K for LBTI down again - OCAM is cooling so this is an unsafe state
  • UT 00:20 - doing an init all - got an UNKNOWN LUCI-2 MOS error - will check on that
    • strain gauges are on again - just like last night
    • zeroed SGs and init - looks OK
  • UT 00:24 - taking test exposures -
    • L1 0001-0002 - 2.51sec dark
    • L2 0001-0002 - 2.51 sec dark
  • UT 00:25 - will check on N30FIeldstop on LUCI-1 (hopefully not as much trouble as yesterday)
    • L1 0003-0004
    • looks good enough, putting blinds back into place
  • UT 00:43 - while waiting, going to check N375 fieldstop on L2
    • L2 0003-0004
    • looks good enough - OBS2 is so sluggish that tweaking it more probably will cause more issues
    • OBS-2 is really damn slow now - bringing up OBS32-bit as a backup just in case - having only two computers for observing is ridiculous
  • UT 00:50 - putting blinds into LUCI-2 - both LUCIs ready for 2nd half.
  • UT 01:10 - LBTI BCU2K SX reconnected

DEC Time On-sky

  • UT 06:40 - LBTI handing over early to DEC time
  • UT 06:41 - reconfiguring to bino LUCIs - authorizing LUCI-1
OSCO HD 19332
  • UT 06:51 - pointing star
  • UT 06:54 - collimation preset
  • UT 06:57 - AO preset to HD 19332 - R=7.46 - 900Hz
    • UT 07:02 - loop failed - optical gain issue
    • UT 07:04 - had to reset a value to get optical gain working
    • UT 07:05 - loop closed
    • FWHM ~3.91-4 pix = 0.058"-0.06"
    • IQ is pretty cruddy
    • L1 0005 0024- NCPA
    • OSCO.png
ESM Gains Test
M52 - Bin 1 R=10.73
  • UT 07:14 - reconfiguring for bino-LUCI mixed mode - putting N30Fieldstop away
  • UT 07:21 - collimation preset
  • UT 07:25 - sending preset
  • SX shows 1" guider, 1.05" DX guider
  • going straight to closed loop, skipping center pupil
  • WFSmag=11.8
  • UT 07:29 - trying optimize gain
  • UT 07:33 - optimize gain done
  • UT 07:34 - taking data
    • tip-tilt 10.8, HO1 10.58
    • L1 0026 - FWHM AO Ref 2.89pix =0.34" , edge of field 3.6pix-4.3pix = 0.42"-0.5"
    • guidestar vanished on SX; 20",-20" - UPDATED IT 8496
    • L1 0029 - FWHM AO Ref 3.pix = 0.36", edge of field 3.6-4.6pix again = 0.42"-0.5"
    • L1 0025-0032- last frame skyframe
    • L2 0005-0012 - last frame skyframe
  • UT 07:45 - manually offseting absolute 0,0 to change gains
    • tip-tilt 3, HO1 0.5
    • L1 0033 - AO Ref star is 2.78pix=0.32"
    • L1 0034- AO Ref star is 2.93pix= 0.34", edges of field 3.8-4.4pix = 0.44"-0.51"
    • L2 0014 - 4.85pix AO Ref star ~0.58"
  • UT 07:52 - changing gains
    • tip-tilt 6, HO1 0.5
    • L1 0035- AORef 2.91pix = 0.34"
    • L1 0036- AO Ref 2.7pix = 0.31"
    • L2 0016 - same star is 4.72pix = 0.56"
    • L1 0037-0038 - sky frame
    • L2 0017-0018 - sky frame
  • UT 08:00 - changing gains
    • tip tilt 6, HO1 3
    • DX guider shows 0.7"
    • L1 0039 - AO Ref 2.69pix = 0.31",
    • L1 0040 -AO Ref 2.69pix = 0.31". 3.7-4.5pix = 0.43"-0.53"
    • L2 0020 -Ref star is 5.3pix = 0.63"
    • L1 stars have ellpticities ~0.05 or better, L2 same stars have ellipticities ~ 0.2
  • UT 08:06
    • tip tilt 10.0, H01 3
    • DX guider shows 0.8"-0.92"
    • L1 0041-AO Ref 2.53pix=0.29", stars at edge=3.5-4pix=0.41"-0.47"
    • L1 0042-AO Ref 2.59pix =0.30", stars at edge 3.3-4pix=0.38"-0.47"
    • L2 Ref star 7.1pix=0.85"
  • L1 0025-0042
  • L2 0005-0022
M52 Bin 4 (R=15.28 but sent as R=16.51)
  • UT 08:12 - Sending preset
  • UT 08:16 - skipping center pupils
  • Seeing blowing up a bit 1.4" on DX guider
  • WFSmag=16.8
  • Not doing optimize gain until the end
  • UT 08:18 - tiptilt 0.3, HO1 0.3
  • UT 08:20 - loops closed, taking data
    • L1 0043- AO Ref 4.28pix=0.5", edge 6pix=0.72"
    • L2 0023-
    • L1 0044- skyframe
    • L2 0024- skyframe
  • UT 08:26 - 0,0 offset changing gains
    • tiptilt 1, HO1 0.3
    • L1 0045- AO Ref 3.68pix=0.43", outer edges 5.1-5.6pix = 0.6"-0.65"
    • L2 0025- Ref star 6.86pix=0.82", outer edges 5.8-6.3pix = 0.69"-0.75"
  • UT 08:29 - changing gains
    • tiptilt 1, HO1 1
    • L1 0046- AO Ref 3.69pix=0.43", outer edges 4.8-5.2pix = 0.56"-0.61"
    • L2 0026- AO Ref star 5.27pix =0.63" outer edges 6.4pix-7.1pix=0.76"-0.81'
  • UT 08:35 changing gains
    • tiptlt 0.3 HO1 1
    • L1 0047- AO Ref star 4.48pix =0.52"
    • L2 0027-AO Ref star = 10.12pix = 1.21"
  • UT 08:38 - optimizing gains
    • tiptilt = 0.84, HO1 =0.45
    • DX guider 1.2"
    • L1 0048- AO Ref 4.68pix =0.55"
    • L2 0028- Ref star 7.40 pix =0.89"

D-Time Target AZ - Target#2 J0313-1800

  • UT 08:48 - parked L1 rotator and at El 84 - putting N30fieldstop into FPU
  • UT 08:50 - sending preset
  • on-axis AO Ref
  • UT 08:59 - seeing is blwoing up while we close the loop
  • UT 09:02 - loop closed starting exposures
    • L1 0049 - looks elongated
    • L1 0050 - loops opened - DO NOT USE THIS IMAGE
    • L1 0051- looks elongated - not sure if this is a binary or tip-tilt went too high
    • L1 0052- turned off NCPA, did not use GOpt - stable - looks like elongation is real and this is a binary - separation looks to be ~0.05"
    • L1 0053-0054 - NCPA off, no GOpt -
    • L1 0055-0056 - turning on GOpt and NCPA - looks the same as NCPA off and no GOpt - so AO system is not introducing elongation
  • UT 09:12 - DIMM 1.2"
    • L1 0049-0065
  • UT 09:15 - DONE

D-Time Target AZ - Target#1 J0313-1802

  • UT 09:16 - sending preset
  • SX guider shows 1"
  • off-axis AO Ref
  • UT 09:23 - turning on NCPA
  • UT 09:24 - loops closed, starting exposures
  • UT 09:26 - FWHM ~0.5" - AO Ref star is 40" - this is kind of the border between AO/ESM - not sure ESM would gain much if anything given brightness of star
  • UT 09:27 - script is done
  • L1 0066-0069

Eng Test full AO w/N375 - Arp 187

  • UT 09:28 - reconfiguring for Bino-LUCIs
  • UT 09:29 - putting N30fieldstop away after parking rotating moving higher elevation to reconfigure
  • UT 09:31 - sending preset
  • UT 09:35 - DX missed guidestar
  • UT 09:39 - user error - B.Rothberg loaded wrong file and accidentally sent a bino full AO - sent new script
  • UT 09:39 - Sat Nov 5 09:39:19.478 2022 left Preset alert for LUCIFER mode ADAPTIVEACE_ACTIVE: AGw command status: AGW1: Requested position is Out of Range
    Sat Nov 5 09:39:19.498 2022 left Preset alert for LUCIFER mode ADAPTIVEACE_ACTIVE: isPosAccessible: probe cannot move to requested position (SFP) (x,y): 153.983, 11.3002 mm.
    Sat Nov 5 09:39:19.515 2022 left Preset alert for LUCIFER mode ADAPTIVEACE_ACTIVE: guidestar list successfully changed but failed moving probe to position.
    Sat Nov 5 09:39:19.553 2022 left Preset command failed for LUCIFER mode ADAPTIVEACE_ACTIVE with guidestars, AO reference, Sound=presetFailed.wav
  • UT 09:44 - V2 of script w/new guidestar - no idea why DX found the guidestar if it was outside of patrol field in SX
  • UT 09:46 - DX guider shows 0.8"-0.9" - Rmag = 13.4, WFSmag = Bin 3 871Hz
  • UT 09:50 - running GOpt again - DX guider 0.6"-0.8" - NCPA on
  • L1 0070- star near galaxy 2.8pix = 0.32, top left edge 4.14pix = 0.48"
  • L2 0029 - 4.85pix = 0.58" - top left edge 6.8pix = 0.82"
  • UT 10:08 - seeing on DX jumped to 1.6"
  • UT 10:10 - L1 0079-0080 - loops opened - got a bunch of skipframes - DO NOT USE
    • L1 0080 - image jumped - double, L1 0076-0077 OK
  • UT 10:16 - resuming, repeating last two dither positions - DX guider 0.9"-1"
  • UT 10:36 - stopping here
    • L1 0070-0091
    • L2 0029-0050

Eng Test Didymos Full AO

  • UT 10:41 - sending collimation script
    • collimating w/1.3" on SX guider - interesting test in mediocre conditions
  • UT 10:44 - sending NS Ace Track script
  • UT 10:50 - DIMM reads 1"
  • UT 10:51 - having trouble closing the loop - resetting the flat
  • UT 10:53 - light is on OCAM - but seems to be sending wrong corrections - might be too extended
  • UT 10:56 - going to resend, cancelling preset, resending preset
  • UT 10:59 - moving bayside stages to find target - was at edge of field in last preset, seems to have been just outside on this one
  • UT 11:00 - D.Miller found it, WFSmag reported 17.4 - didn't chage to clear filter
  • UT 11:03 - putting clear filter in to see if more flux lands on OCAM
    • went from 17.2 to 16.7 - half mag improvement!
    • 227Hz
  • UT 11:05 - optimize gain
  • UT 11:06 - taking data - got weird error on LUCI
    • 2022.11.05 11:06:40 s | error high | luci.luc | Luci ONE | ONE_GEIRSServer | RMIGEIRSServiceImpl.java#parseFileSystem(1044) | GEIRS received response with error: ERROR 49 Command 'cd /dataA/readout1/DATA/20221105/cubes' returned errorcode = 49: (E_dnotf=49) cannot cd to CAMPATH (image directory)
    • L1 0092-0093 - off-centered because bayside stages were moved
  • UT 11:12 - re-aligning N30fieldstop
    • DET Y -10" put object in field, moving it -15" DET Y should center it -, then weill absorb, do an absolute 0,0,take exposure
    • L1 0094-0104 - did 10NDITS x2.51sec test exposure to see if this worked - it worked! but i'm dyslexic - was supposed to be 10DIT
  • UT 11:16 - moving to 2nd line in script and going
  • UT 11:41 - DIMM 1"
    • L1 0105-0124 (0118 is 3DIT x 20 NDIT instead of 20 DIT x 3NDIT)
    • L1 0124 - has interference in it - lines
PSF Star SDSS J080336.58+053430.6 - AO star
  • UT 11:51 - sending preset
  • Guider shows 0.92"
  • WFSmag = 17.4
  • 157Hz
    • L1 0125-0130

D-Time Target AZ - Target#4 J0313-1802

  • UT 12:08 - sending preset
  • DIMM 1.2"
  • UT 12:18 - loops closed, starting exposures
  • UT 12:19 - looks like its saturated at K-band
    • L1 0131-0132 - DO NOT USE - SATURATED
  • UT 12:20 - putting in Br Gamma filter
  • UT 12:23 - no evidence of a binary - FWHM ~ 4pix ~0.06"
  • last exposure manually offseting to -4, 4 to replace lost exposure
  • UT 12:25 - complated
    • L1 0133-0136

OSCO - HD 61059

  • UT 12:31 - sending preset
  • UT 12:34 - accidentally put blind filters in - thanks to sluggish OBS2
  • UT 12:37 - resending preset
  • Rmag = 7.49, WFSmag=8.7, which matches the offset from the OT
  • Getting 127Hz
  • UT 12:26 - turning down gains
  • UT 12:50 2022.11.05 12:49:40 s | error medium | luci.luc | LBT | TelescopeServer | CommandThread.java#run(469) | RunAO failed
    • No corrections being sent, above was the error on the LUCI message daemon
    • Corrections started up again with optimize gain
    • now getting skipframes
  • UT 12:54 - LUCI timed out, taking data manually
    • L1 0137-0155
  • UT 1300 - Closing DOME

Closed Dome Cals

N30 LUCI-1 Cals
  • L1 0156-0165 - H-band flats
  • L1 0166-0175 - BrGamma flats
  • L1 0176-0185 - K-band flats
N375 Bino Cals
  • L1 0186-0195 BrGamma ; L2 0051-0060 - BrGamma
  • L1 0196-0205 Ks ; L2 0061-0070 Ks
  • L1 0206-0215 K ; L2 0071-0080 K
Topic revision: r7 - 05 Nov 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback